OpenStudy (anonymous):

The area of a triangle is 1,440 cm2. The base of the triangle is 5 times the height. What is the height of the triangle? 12 cm 24 cm 46 cm 60 cm

OpenStudy (anonymous):

46?

OpenStudy (ghostgate):

To find the area of a triangle you would use this formula: bh/2 = Area. If the base is 5 times the height, then 46 wouldn't be the answer. That would be: 46 x 230 which equals 10,580 / 2 = 5,290 which isn't 1,440. However if it was 24 we get this: 24 x 5 = 120 24 x 120 = 2,880 / 2 = 1,440. Which is our answer. So the answer is: B) 24 cm